The sum of Two Numbers Is 30 and There difference Is 12. Find The Two numbers
Answer by amarjeeth123(569)

Let one number be x.
Then the other number is (30-x) because the sum of the numbers is 30.
x-(30-x)=12
x+x=30+12
2x=42
x=21
The numbers are 21 and 9 respectively.